### v2

# Astronomical Cardiology: A Search For Heartbeat Stars Using Gaia and TESS

Abstract
Heartbeat stars are a subclass of binary stars with short periods, high eccentricities, and phase-folded light curves that resemble an electrocardiogram. We start from the Gaia catalogs of spectroscopic binaries and use TESS photometry to identify 112 new heartbeat star systems. We fit their phase-folded light curves with an analytic model to measure their orbital periods, eccentricities, inclinations, and arguments of periastron. We then compare these orbital parameters to the Gaia spectroscopic orbital solution. Our periods and eccentricities are consistent with the Gaia solutions for 85 of the single-line spectroscopic binaries but only 20 of the double-line spectroscopic binaries. For the two double-line spectroscopic binary heartbeat stars with consistent orbits, we combine the TESS phase-folded light curve and the Gaia velocity semi-amplitudes to measure the stellar masses and radii with PHOEBE. In a statistical analysis of the heartbeat star population, we find that non-giant heartbeat stars have evolved off the main sequence and that the fraction of the systems that are heartbeat stars rises rapidly with effective temperature.
keywords:
binaries: eclipsing – binaries: spectroscopic1. Introduction
Heartbeat stars (HB) are a relatively rare subclass of detached binary stars with short periods ( days) and high eccentricities (). In these binaries, the stars are tidally deformed near pericenter, causing periodic brightness variations. HBs are easily identifiable because their phase-folded light curves resemble an electrocardiogram (Thompson et al., 2012), and a simple analytic model (Kumar et al., 1995) can be used to estimate the properties of the orbit. The tides can also produce tidally excited oscillations (TEOs) of the star that are observed between pericenter passages.
Since the amplitude of the photometric variations in HBs is typically small, it was only with the Kepler mission that HBs were detected (Welsh et al., 2011). KOI-54 was one of the first HBs from Kepler, and it also had TEOs (Welsh et al., 2011; Fuller & Lai, 2012; Burkart et al., 2012; O’Leary & Burkart, 2014). However, HBs would not have their name until Thompson et al. (2012) applied the analytic Kumar et al. (1995) model to fit the Kepler phase-folded light curves of 17 new HB systems. Hundreds of HBs were identified in Kepler (Hambleton et al., 2013; Kirk et al., 2016; Shporer et al., 2016). These systems are all on the main sequence (MS) and have periods of days. The Optical Gravitational Lensing Experiment (OGLE) has discovered nearly 1000 HBs across the upper MS and giant branch (Wrona et al., 2022a, b).
The Transiting Exoplanet Survey Satellite (TESS, Ricker et al., 2015) provides another opportunity to discover short period ( days) HBs. TESS observes the sky in twenty-six sectors which are continuously observed for roughly twenty-seven days. Like in Kepler, early detections of HBs in TESS came as a byproduct of planetary searches (Wheeler & Kipping, 2019). TESS photometry has also been used to identify or characterize HB variability in known binaries like MACHO 80.7443.1718 (Jayasinghe et al., 2019, 2021; Kołaczek-Szymański et al., 2022), KIC 5006817 (Merc et al., 2021), V680 Mon (Paunzen et al., 2021), WR 21a (Barbá et al., 2022), and FX UMa (Wang et al., 2023). HBs are also discovered serendipitously in various TESS star samples (Murphy et al., 2020; Kochukhov et al., 2021; Sharma et al., 2022; Martín-Ravelo et al., 2024).
There have also been systematic surveys for HBs in TESS. By analyzing the first dozen TESS sectors, Kołaczek-Szymański et al. (2021) found 20 new HB systems through the visual inspection of the phase-folded light curves of spectroscopic binaries. Li et al. (2024a, b) doubled the number of HBs found by deliberate searches in TESS by cross-matching and visually inspecting the TESS light curves of the stars in the PPM catalog (Röser & Bastian, 1993; Röser et al., 1994; Nesterov et al., 1995). Recently, Solanki et al. (2025) identified 180 HBs using the TESS full-frame images by applying convolutional neural networks to TESS phase-folded light curves, followed by a visual inspection to confirm the candidates.
In contrast to previous HB searches, which start from large samples of phase-folded light curves, here we start from the Gaia catalogs of double-line (SB2) and single-line (SB1) spectroscopic binaries. This has the advantage of starting from a sample of known binaries, and we can then use the SB2 HB systems to estimate the masses and radii of the stars. In Section §2, we discuss the Gaia data and how we model and search their phase-folded TESS light curves for HBs. Sections §2.1 and §2.2 describe our search strategies for HBs in the Gaia SB2 and SB1 samples, respectively. In Section §3, we compare the orbital parameters we measure from the TESS phase-folded light curves to the Gaia orbital solutions. We also model the SB2 HBs with consistent Gaia/HB orbital parameters to determine the masses and radii of the systems. Finally, Section §4 discusses the statistics of HBs and future steps.
2. Heartbeat Search
We start from the 181,529 SB1s and 5,376 SB2s (Gaia Collaboration et al., 2023) in Gaia DR3. We used phase-folded light curves for sectors 1-79 from the TESS Quick-Look Pipeline (QLP, Huang et al., 2020) for all of the Gaia SB1s and SB2s. The typical Gaia spectroscopic binary has four TESS sectors of observations.
2.1 SB2 Modeling
| TIC | Type | Score | incl () | () | P (days) | Gaia P (days) | P frac diff | e | Gaia e | Gaia e err | Match | Eclipses | TEOs | |
| 1598625 | SB1 | 0.442 | 36.130 | 232.130 | 3.252 | 3.241 | 0.004 | 0.330 | 0.404 | 0.096 | ✓ | ✗ | ✗ | |
| 9175061 | SB1 | 0.476 | 55.720 | 141.410 | 2.900 | 2.900 | 0.000 | 0.160 | 0.147 | 0.034 | ✓ | ✗ | ✓ | 2 |
| 11578775 | SB1 | 0.236 | 48.110 | 115.970 | 5.327 | 5.331 | 0.001 | 0.180 | 0.180 | 0.104 | ✓ | ✗ | ✗ | |
| 13933268 | SB1 | 0.374 | 51.480 | 308.690 | 4.129 | 4.126 | 0.001 | 0.110 | 0.127 | 0.019 | ✓ | ✗ | ✗ | |
| 15257961 | SB1 | 0.468 | 38.260 | 360.000 | 7.748 | 7.763 | 0.002 | 0.110 | 0.137 | 0.053 | ✓ | ✗ | ✗ | |
| 22567490 | SB1 | 0.320 | 56.800 | 61.110 | 6.567 | 6.554 | 0.002 | 0.550 | 0.250 | 0.034 | ✗ | ✓ | ✗ | |
| 26412885 | SB1 | 0.474 | 75.900 | 85.750 | 3.490 | 3.490 | 0.000 | 0.160 | 0.108 | 0.045 | ✓ | ✗ | ✗ | |
| 26689977 | SB1 | 0.262 | 40.260 | 301.160 | 3.692 | 3.697 | 0.001 | 0.110 | 0.125 | 0.053 | ✓ | ✗ | ✗ | |
| 28543727 | SB1 | 0.562 | 90.000 | 26.480 | 2.576 | 2.579 | 0.001 | 0.190 | 0.130 | 0.022 | ✗ | ✗ | ✗ | |
| 29451470 | SB1 | 0.497 | 65.910 | 343.170 | 3.011 | 3.260 | 0.076 | 0.110 | 0.102 | 0.043 | ✓ | ✗ | ✓ | 2 |
| 29518898 | SB1 | 0.376 | 90.000 | 306.270 | 4.649 | 4.628 | 0.004 | 0.130 | 0.163 | 0.074 | ✓ | ✗ | ✗ | |
| 31096993 | SB1 | 0.219 | 33.590 | 356.980 | 6.019 | 6.018 | 0.000 | 0.160 | 0.191 | 0.031 | ✓ | ✗ | ✗ | |
| 42430804 | SB1 | 0.239 | 45.180 | 216.590 | 8.998 | 0.403 | 21.314 | 0.560 | 0.168 | 0.060 | ✗ | ✗ | ✗ | |
| 42741990 | SB1 | 0.336 | 37.350 | 319.100 | 5.785 | 5.817 | 0.005 | 0.150 | 0.094 | 0.102 | ✓ | ✗ | ✗ | |
| 44787510 | SB1 | 0.720 | 22.360 | 360.000 | 3.553 | 1.860 | 0.910 | 0.470 | 0.096 | 0.019 | ✗ | ✗ | ✗ | |
| 61160125 | SB1 | 0.138 | 48.110 | 345.120 | 5.960 | 5.908 | 0.009 | 0.220 | 0.396 | 0.100 | ✗ | ✗ | ✗ | |
| 63548665 | SB1 | 0.286 | 50.370 | 298.410 | 4.770 | 1.124 | 3.243 | 0.190 | 0.057 | 0.047 | ✗ | ✗ | ✗ | |
| 65315466 | SB1 | 0.554 | 90.000 | 226.240 | 3.196 | 3.196 | 0.000 | 0.190 | 0.232 | 0.086 | ✓ | ✗ | ✗ | |
| 73486038 | SB1 | 0.771 | 57.060 | 85.460 | 6.445 | 6.451 | 0.001 | 0.260 | 0.302 | 0.048 | ✓ | ✗ | ✗ |
Since there are only 5,376 Gaia SB2 targets, we identified HB candidates through visual inspection of the TESS phase-folded light curves. For all targets, we run a Lomb-Scargle periodogram (LS, Lomb, 1976; Scargle, 1982) on each TESS sector independently. We visually inspected the phase-folded light curves and identified 10 HB candidates. We then use phase dispersion minimization (PDM, Stellingwerf, 1978) to refine the orbital period. Before modeling the phase-folded light curve, we select a single TESS sector for each target and remove long-term trends in the data, if necessary. For example, TIC 265056364 has a continuous decrease in flux over the course of the TESS sector, which could be due to variable blended light or systematic trends. We remove these signals with a linear fit.
We fit the phase-folded light curves using the Kumar et al. (1995) analytic model. The flux is modeled as
where sets the amplitude, is the mean flux, is the argument of periastron, is the inclination, and is the orbital eccentricity. The true anomaly is
and the eccentric anomaly
is determined by the period, and the epoch of periastron, . The shape of the phase-folded light curve varies significantly with the orbital configuration, as shown in Fig. 1.
We used Markov Chain Monte Carlo (MCMC) methods as implemented by Foreman-Mackey et al. (2013) to sample over the parameters in the model. We ran the MCMC models with fourteen walkers, ten thousand iterations, and a burn-in of one hundred iterations. Fig. 2 shows the model fits for nine of the SB2 HBs and Table 1 reports the median estimates for the argument of periastron, inclination, eccentricity, and period.
2.2 SB1 Modeling
Since there are 181,529 Gaia SB1 orbital solutions, we performed a semi-automated search rather than relying only on visual inspection. We only analyzed the 43,599 systems with Gaia orbital periods less than 13 days so that there would be at least two binary orbits in a TESS sector. The TESS light curves often have features at the start and end of each sector leading to significant numbers of false positives for longer periods. For each light curve sector, we used a LS periodogram to identify periodic variables. We then bin the phase-folded light curves into one hour time-bins, and fit the Kumar et al. (1995) analytic model (Eq. 1) to the phase-folded light curve using the Trust Region Reflective algorithm. Since the LS periodogram often returns a simple fraction of the orbital period, we do this for , , and where is the period corresponding to the maximum power in the periodogram.
For each target, we select the sector and period combination that minimizes the reduced of the Kumar model fit. We then compare this to the for a linear fit to the phase-folded light curves. Fig. 3 shows the distribution of and orbital eccentricities for the 43,599 systems. We use the HBs identified in the SB2 sample to define the selection region ( and ) shown in Fig. 3. This left 7,314 binaries, which we then visually inspected to select 102 HB candidates. Fig. 4 shows examples of the phase-folded light curves with their Kumar model fits, and Table 1 reports the orbital parameters.
Some of our HB candidates also have primary or secondary eclipses, such as TIC 147307851, TIC 143634957, and TIC 22567490 in Fig. 4 (they are flagged in Table 1). Since the Kumar model only accounts for tidal distortions, we manually mask the eclipses when necessary and re-fit the phase-folded light curve. In some cases, the revised fit pushes the eccentricity below the 0.15 cutoff used for selection, but we keep these systems in our final sample.
2.3 The Final Sample
Fig. 5 shows the orbital periods and eccentricities of the 112 HBs detected in the Gaia SB1 and SB2 samples. Of these 112 HBs, four were found by Solanki et al. (2025) (TIC 252588526, TIC 271554988, TIC 344586348, and TIC 370209445). We found no matches to the other comparison samples used below. The orbital periods range from 1.51 to 12.17 days, with a median of 4.36 days. The eccentricities range from 0.10 to 0.57 (TIC 42430804), with a median of 0.20. All of our targets fall below the period-eccentricity envelope expected from tidal circularization (Mazeh, 2008). Fig. 5 also shows the HBs from Wrona et al. (2022a, b), Li et al. (2023), and Solanki et al. (2025). As compared to the HBs identified in Wrona et al. (2022a, b), our systems have much shorter periods than their median value of 265.3 days. The median eccentricities are similar (0.24 for OGLE). Our period cut will exclude giants while the OGLE magnitude limits favour giants given the distance to the Magellanic Clouds.
Fig. 6 shows the stars on an extinction-corrected Gaia CMD. We use distances from Bailer-Jones et al. (2021) and extinctions from the mwdust three-dimensional dust map (Bovy et al., 2016; Drimmel et al., 2003; Marshall et al., 2006; Green et al., 2019). The candidates have good parallaxes () and none are highly extincted ( mag), so their CMD locations should be reliable. As compared to the full sample of Gaia spectroscopic binaries, the HBs tend to be higher on the MS (median ). Fig. 6 also includes 478 HBs from OGLE (Wrona et al., 2022a, b), 160 from Kepler (Shporer et al., 2016; Li et al., 2023; Dimitrov et al., 2017; Kirk et al., 2016), and 179 from TESS (Kołaczek-Szymański et al., 2021; Solanki et al., 2025).
We used periodograms of the residuals of the Kumar et al. (1995) models, to search for TEOs. We list the systems with periodogram powers larger than a false alarm level of at a simple harmonic of the orbit in Table 1. Ten of the 102 SB1 HB systems show TEOs, Fig. 7 shows three examples. Similar to Li et al. (2024b), our TEOs appear at low orbital harmonics with a median value of and a maximum at for TIC 173561516. The TEOs can be used to study the internal structure of these stars.
3. Gaia Comparison
Fig. 8 compares the Gaia spectroscopic orbit solutions and HB fit orbital periods, eccentricities, and arguments of periastron. We were particularly interested in the SB2s because agreement between the orbital parameters means that the Gaia velocity semi-amplitudes can be combined with the photometric variability to measure masses and radii (Rowan et al., 2023). Unfortunately, only 2 of the 10 SB2 HBs have orbital periods and eccentricities that are in good agreement. It is difficult to pinpoint why an individual Gaia orbital solution is incorrect, since the epoch RVs are not included in Gaia DR3, but it is likely due to the large number of degenerate periods that are possible when fitting sparse RVs measured over a long time baseline to short period orbits. Furthermore, most of the SB2 HBs are B/A stars on the upper MS (Fig. 6), and the 846–870nm wavelength range of the Gaia RVS instrument was primarily designed for cool stars (Cropper et al., 2018). This is one reason that the RVs for hot stars only became available with DR3 (Blomme et al., 2023).
The SB1 HBs have considerably more accurate Gaia orbits, with 85 of the orbital periods and eccentricities agreeing with the light curve solution. Bashi et al. (2022) introduced a “score” statistic ranging from to quantify the accuracy of Gaia SB1 orbital solutions. Of the 87 matching SB1 HB systems, 73 have scores below the cutoff recommended by Bashi et al. (2022) for the selection of a “clean” SB1 sample. Despite this, the score statistic seems to be a limited indicator of the Gaia orbit quality, as almost all of the targets with incorrect Gaia , , or also have .
For the two SB2 HB systems with correct Gaia periods and eccentricities we used the PHysics Of Eclipsing BinariEs (PHOEBE) package, a Python package for modeling EB systems (Prša & Zwitter, 2005; Prša et al., 2016; Conroy et al., 2020), to determine the stellar masses and radii. We use the period and orbital parameters from the HB models as the initial conditions. We use the Gaia and velocity amplitudes to constrain the mass ratio
and the projected semi-minor axis of the companion
Only the uncertainties in and are important. We initialize the PHOEBE MCMC walkers near the values from the initial conditions, use Gaussian priors on and , and then run the MCMC minimization with twenty walkers and five hundred iterations.
Fig. 9 shows the posteriors for TIC 98552498. The mass and radius of the primary are (2.61 0.27) and (1.41 0.40) , while the mass and radius of the secondary are (2.53 0.29) and (2.85 0.11) . Fig. 10 shows the posteriors for TIC 76094846. The mass and radius of the primary are (8.19 1.19) and (5.14 0.18) , while the mass and radius of the secondary are (8.11 1.39) and (4.97 0.47) . Both systems are consistent with the masses and radii of massive stars measured from detached eclipsing binaries (Torres et al., 2010).
4. Discussion
We used the TESS light curves of Gaia spectroscopic binaries to identify one of the largest samples of HBs, finding 112 HBs among the 186,905 spectroscopic binaries in the Gaia DR3 catalog, of these, 18 show eclipses, and 10 have TEOs. Out of the ten SB2 HBs, only two have orbital parameters in good agreement with the Gaia values. The problem likely arises because of the sparse Gaia RVs, the short periods, and that most of the systems are hot stars (Fig. 6). For the SB1 HBs, 85 of the orbits agreed with the Gaia estimates.
One of the original goals of this project was to find the SB2 HB systems and then model them to determine the component masses and radii. The problems with the SB2 solutions meant we could attempt this for only two systems. TIC 98552498 consists of a primary with mass (2.61 0.27) and radius (1.41 0.40) , and a secondary with mass (2.53 0.29) and radius (2.85 0.11) . TIC 76094846 consists of a primary with mass (8.19 1.19) and radius (5.14 0.18) , and a secondary with mass (8.11 1.39) and radius (4.97 0.47) . To analyze the remaining SB2 HBs, we require RV measurements from independent sources or the actual Gaia RV measurements (which will be released in DR4).
There are two striking features of the distribution of HBs along the MS of the CMD shown in Fig. 6. First, the HBs seem to be more luminous at fixed color than their parent samples. Second, there seem to be many fewer HBs on the lower MS. To quantify these issues, we first applied a cut removing stars with to eliminate most of the giant branch (Fig. 11). This is imperfect for stars redder than roughly mag, particularly for the SB1 systems. Fig. 11 also shows the 0.5 Gyr MIST isochrone (the "single" isochrone) and the same isochrone shifted to be 0.75 mag brighter (the "binary" isochrone), the shift expected for a binary of stellar twins. These are just to guide the eye since they are not a good match to the observed upper MS.
To examine the first point, we computed the median magnitudes of the samples in colour bins with a width of mag. The medians of the full (or period trimmed) SB1 and SB2 samples behave as expected on the upper MS. They roughly track the isochrones, and the SB2 magnitudes tend to be a little more luminous since having a more luminous companion will favor being able to obtain an SB2 solution. On the lower MS, the scatter of more luminous stars above the MS but below our cut against giants affects the SB1 sample more strongly than the SB2 sample because having an evolved primary makes measuring the velocity of the secondary less likely. The median absolute magnitudes of the HBs (SB1, SB2, or all HBs) are then more luminous than the SB2 sample. This strongly indicates that HBs are not just binaries, but binaries in which the primary has started to evolve. Following the logic in MacLeod & Loeb (2025), this is a natural consequence of HB amplitudes being higher as stellar evolutionary time scales become shorter.
To examine the second point, we computed the fraction of HBs as a function of color. There are two reasons the HB fraction might drop as we examine stars lower on the MS. First, the redder, lower mass stars have longer and longer evolutionary time scales, which will lead to lower amplitudes (MacLeod & Loeb, 2025). Second, stars below the Kraft break ( K or for the isochrone in Fig. 11, Kraft, 1967; Schatzman, 1962; Beyer & White, 2024) have convective envelopes and so should have shorter damping time scales. Both of these effects will lead to lower HB amplitudes. We have some ability to distinguish the two issues because the highest mass stars below the Kraft break (Kraft, 1967; Schatzman, 1962; Beyer & White, 2024) can be old enough to have started to evolve off the MS. We computed the fractions of SB1 and SB2 HBs as a function of color again, but along lines parallel to the one labeled "Kraft break" in Fig. 11 in mag bins. We report the median fractions and 90 confidence uncertainties or upper limits.
The results shown in Fig. 12 are striking. The fraction of HBs drops very rapidly for redder, cooler stars starting above the Kraft break (Kraft, 1967; Schatzman, 1962; Beyer & White, 2024) and then may be flattening for still cooler stars. The SB2 fractions and limits are in agreement with those for the SB1s. The absolute normalizations of the SB1 fractions change when we use the full SB1 sample for the normalization of the fraction, rather than just those with periods days, but the trends are unchanged. On the upper MS, HBs are not rare - over 1 of Gaia SB1 systems bluer than mag are HBs.
Starting from a sample of binaries can simply be viewed as a means of increasing the efficiency of HB searches, particularly since many catalogs are dominated by shorter period systems. There are significant systematic effects in the catalogs which would need to be taken into account for any statistics as a function of period. This can be done more broadly by selecting Gaia or APOGEE (e.g., Abdurro’uf et al., 2022) stars with significant velocity scatters (e.g., Badenes et al., 2018) rather than requiring an orbit solution. It would be good to find a solution to the period search limitations of TESS. Evolved HBs with longer periods may be detectable in the upcoming TESS Cycle 8, which includes longer 54 day sectors and “rolled” sectors where some parts of the sky will be observed over consecutive sectors. It might also be possible to search for longer periods in TESS and then use heavily binned All-Sky Automated Survey for Supernovae (ASAS-SN, Shappee et al., 2014; Kochanek et al., 2017) light curves to verify the periods and characterize the HB signal. While directly searching for such weak signals in ASAS-SN might well be problematic, Hon et al. (2025) have shown that binned ASAS-SN light curves can detect very weak signals.
It is clear that there is considerable physics in characterizing and understanding the distribution of HBs as a function of orbital and stellar properties. What we qualitatively observe can largely be explained as the competition between evolutionary and dissipation rates, as discussed by MacLeod & Loeb (2025). High mass stars evolve faster and have less dissipation, leading to larger numbers of HBs. There are, however, a large number of variables (mass, evolutionary state, orbit, etc.), which means that significantly larger samples will be needed to characterize the physical parameter space of HBs fully.
Acknowledgments
JC thanks Tharindu Jayasinghe and the Time Domain Research Group for their guidance, advice, and support. JC also thanks Collin Christy for his assistance through the years. DMR acknowledges support from the OSU Presidential Fellowship. CSK, KZS, and DMR are supported by NSF grants AST-2307385 and 2407206. ASAS-SN is funded by Gordon and Betty Moore Foundation grants GBMF5490 and GBMF10501 and the Alfred P. Sloan Foundation grant G-2021-14192. Thanks to Dr. J. Tayar for a communication on an earlier version of this manuscript.
Our work presents results from the European Space Agency space mission Gaia. Gaia data are being processed by the Gaia Data Processing and Analysis Consortium (DPAC). Funding for the DPAC is provided by national institutions, in particular the institutions participating in the Gaia MultiLateral Agreement. This paper also includes data collected by the TESS mission. Funding for the TESS mission is provided by the NASA’s Science Mission Directorate.